OpenStudy (anonymous):

A skyscraper is 396 meters tall. At a certain time of day, it casts a shadow that is 332 meters long. At what angle is the sun above the horizon at that time?

OpenStudy (anonymous):

bump

OpenStudy (anonymous):

Scarlett is trying to find the height of a dam. She stands 90 meters away from the dam and records the angle of elevation to the top of the dam to be 26º.

OpenStudy (anonymous):

the scarlett one is my question

OpenStudy (anonymous):

i copied and posted the wrong question

OpenStudy (igreen):

Okay, so we have a right triangle..

OpenStudy (igreen):

We have the adjacent and we want to find the opposite. Use tangent. \(\sf tangent = \dfrac{Opposite}{Adjacent}\) \(\sf tan(20^o) = \dfrac{x}{90}\) Plug in tan 20 degrees into your calculator..what do you get?

OpenStudy (igreen):

Or put it in google.

OpenStudy (anonymous):

0

OpenStudy (anonymous):

or 0.22

OpenStudy (anonymous):

???

OpenStudy (anonymous):

bump